# Reinforcement Unlearning

### v2

@@ Abstract.
Machine unlearning refers to the process of mitigating the influence of specific training data on machine learning models based on removal requests from data owners. However, one important area that has been largely overlooked in the research of unlearning is reinforcement learning. Reinforcement learning focuses on training an agent to make optimal decisions within an environment to maximize its cumulative rewards. During the training, the agent tends to memorize the features of the environment, which raises a significant concern about privacy. As per data protection regulations, the owner of the environment holds the right to revoke access to the agent’s training data, thus necessitating the development of a novel and pressing research field, termed reinforcement unlearning. Reinforcement unlearning focuses on revoking entire environments rather than individual data samples. This unique characteristic presents three distinct challenges: 1) how to propose unlearning schemes for environments; 2) how to avoid degrading the agent’s performance in remaining environments; and 3) how to evaluate the effectiveness of unlearning. To tackle these challenges, we propose two reinforcement unlearning methods. The first method is based on decremental reinforcement learning, which aims to erase the agent’s previously acquired knowledge gradually. The second method leverages environment poisoning attacks, which encourage the agent to learn new, albeit incorrect, knowledge to remove the unlearning environment. Particularly, to tackle the third challenge, we introduce the concept of “environment inference” to evaluate the unlearning outcomes.
1. Introduction
@@
Machine learning relies on the acquisition of vast amounts of data, which may encompass sensitive information of individuals. To safeguard the privacy of individual users, data protection regulations have been proposed, e.g., the General Data Protection Regulation (GDPR) (GDPR, 2016), which empowers users to request the removal of their data. It is imperative for model owners to adhere to users’ requests by removing revoked data from their datasets and ensuring that any influence these revoked data may have on the model is eliminated. This process is referred to as machine unlearning (Cao and Yang, 2015; Bourtoule et al., 2021).
@@
While significant progress has been made in conventional machine unlearning (Bourtoule et al., 2021; Guo et al., 2020; Zhang et al., 2022), one area that remains an unfilled gap for unlearning is reinforcement learning (RL). RL is an essential research field in machine learning due to its ability to address complex decision-making problems in dynamic environments (Wang et al., 2022a). In RL, the primary objective is to train an intelligent entity, known as an agent, to interact with the environment through a specific policy. This policy guides its actions based on the current state. With each action taken, the agent receives a reward and consequently updates its state, creating an experience sample used to update its policy. The ultimate aim of the agent is to learn an optimal policy that maximizes its cumulative rewards over time.
@@
However, in the course of RL, agents tend to memorize features of their environments, raising substantial security concerns. Consider an RL agent designed for providing navigation guidance through real-time data from Google Maps. During its training, the agent learns from a dynamic environment using Google Street Views for its photographic content (Mirowski et al., 2018). However, privacy issues may arise when the agent inadvertently learns and stores sensitive information, such as the locations of restricted areas. Without a forgetting mechanism, this poses a significant privacy risk, potentially compromising the anonymity of individuals and organizations.
In addition, the agent’s ability to forget is also crucial for ensuring the security of autonomous vehicles. A security challenge arises from malicious entities that exploit the mapping system by registering fraudulent businesses on Google Maps (Huang et al., 2017). Their aim is to redirect organic search traffic away from legitimate establishments and steer it towards profit-driven scams. When training an autonomous driving agent, utilizing maps tainted by such fraudulent entries can lead to suboptimal user experiences or even result in fatal consequences. Thus, the need for the RL agent to forget such sensitive or incorrect information gives rise to a novel research field, which we term as reinforcement unlearning.
@@
Additionally, reinforcement unlearning is distinct from privacy-preserving RL (Vietri et al., 2020; Garcelon et al., 2021). Reinforcement unlearning aims to selectively erase learned knowledge from the agent’s memory, ensuring the privacy of environment owners, while privacy-preserving RL focuses on preserving the agent’s personal information. In essence, reinforcement unlearning presents three specific challenges.
  • How can we unlearn an environment from the agent’s policy? In machine unlearning, a data owner can specify which data samples should be removed. However, in reinforcement unlearning, the environment owner cannot access the experience samples. The difficulty arises as these samples are dynamically accumulated during the agent’s interactions with the environment, and samples are managed by the agent. Thus, the key challenge lies in effectively associating the environment that needs to be unlearned with the corresponding experience samples.
  • How can we prevent a degradation in the agent’s performance after unlearning? In conventional machine unlearning, removing a sample may lead to a decrease of performance. It is more challenging in reinforcement unlearning as unlearning an environment requires forgetting a significant number of experience samples. Ensuring that the agent maintains its performance on the retained environments becomes a considerable challenge.
  • How can we evaluate the effectiveness of reinforcement unlearning? In machine unlearning, one commonly used evaluation is using membership inference attack (Shokri et al., 2017; Ye et al., 2022; Salem et al., 2023) to assess if the model has discarded the revoked data. However, this methodology cannot be directly applied to reinforcement unlearning, as the environment owner cannot specify which samples should be unlearned. This poses a challenge in evaluating the effectiveness of reinforcement unlearning and in determining if the agent has been over- or under-unlearned.
To address these challenges, we propose two distinct unlearning methods: decremental reinforcement learning and environment poisoning. Decremental reinforcement learning involves deliberately erasing an agent’s learned knowledge about a specific environment. This method finds practical applications in scenarios where certain environments become obsolete or need to be forgotten due to privacy concerns. Environment poisoning-based method aims to create poisoning experience samples by modifying the unlearning environment. This method ensures that the agent’s performance in other environments remains unaffected, eliminating any negative impact on its overall capabilities. This method finds application in situations where attacks or misinformation may be present. Both methods enable an agent to unlearn specific environments while maintaining its performance in others, thereby tackling the first two challenges. To tackle the third challenge, we utilize environment inference to infer an agent’s training environment by observing its behavior. If the inference result after unlearning shows a substantial degradation compared to the result before unlearning, it is indicative that the agent has effectively unlearned that environment.
@@
  • Reinforcement unlearning exposes an impactful vulnerability of RL – the risk of exposing the privacy of the environment owner. This vulnerability can disclose sensitive information about the environment owner’s preferences and intentions. We introduce two innovative reinforcement unlearning methods: decremental RL-based and environment poisoning-based approaches.
  • With limited prior research in reinforcement unlearning, to confirm the unlearning results, we introduce a novel evaluation approach, “environment inference”. By visualizing the unlearning results, this approach provides an intuitive and effective means of measuring the efficacy of unlearning techniques.
2. Preliminaries
Reinforcement Learning.  The primary objective of RL is to learn an optimal policy for the agent, enabling it to maximize the total accumulated reward and accomplish the task optimally. In the context of deep reinforcement learning (DRL), the policy is typically represented by a deep neural network.
Formally, a learning environment is commonly formulated by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},r\rangle$ (Mnih and et al., 2015). Here, $\mathcal{S}$ and $\mathcal{A}$ denote the state and action sets, respectively, while $\mathcal{T}$ represents the transition function, and $r$ represents the reward function. At each time step $t$, the agent, given the current environmental state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$ based on its policy $\pi(s_{t},a_{t})$. This action causes a transition in the environment from state $s_{t}$ to $s_{t+1}$ according to the transition function: $\mathcal{T}(s_{t+1}|s_{t},a_{t})$. The agent then receives a reward $r_{t}(s_{t},a_{t})$, along with the next state $s_{t+1}$. This tuple of information, denoted as $(s_{t},a_{t},r_{t}(s_{t},a_{t}),s_{t+1})$, is collected by the agent as an experience sample utilized to update its policy $\pi$. Typically, the policy $\pi$ is implemented using a Q-function: $Q(s,a)$, estimating the accumulated reward the agent will attain in state $s$ by taking action $a$. Formally, the Q-function is defined as:
(1) $Q_{\pi}(s,a)=\mathbb{E}_{\pi}[\sum^{\infty}_{i=1}\gamma^{i}\cdot r(s_{i},a_{i})|s_{i}=s,a_{i}=a],$
@@ 2. Preliminaries
In deep reinforcement learning, a neural network is employed to approximate the Q-function, denoted as $Q(s,a;\theta)$, where $\theta$ represents the weights of the neural network. The neural network takes the state $s$ as input and produces a vector of Q-values as output, with each Q-value corresponding to an action $a$. To learn the optimal values of $Q(s,a;\theta)$, the weights $\theta$ are updated using a mean squared error loss function $\mathcal{L}(\theta)$.
(2) $\mathcal{L}=\frac{1}{|B|}\sum_{e\in B}[(r(s_{t},a_{t})+\gamma\max_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta)-Q(s_{t},a_{t};\theta))^{2}],$
@@
Machine Unlearning.  Machine unlearning focuses on removing specific data samples or learned knowledge from a trained model. Unlike model training, unlearning aims to erase the impact of certain data samples on the model’s behavior. A straightforward unlearning method involves removing the revoked data and retraining the model from scratch. However, this approach is computationally challenging. To improve computational efficiency, several machine unlearning methods have been proposed. SISA (Sharded, Isolated, Sliced, and Aggregated) is one of the most prevalent methods (Bourtoule et al., 2021), aiming to divide the training set into disjoint shards and train each shard model separately. When a revoke request is received, only the corresponding shard model is retrained.
When combining machine unlearning with reinforcement learning, a novel concept called reinforcement unlearning is introduced. This concept revolves around the selective removal or modification of the agent’s acquired knowledge within the context of reinforcement learning, offering unique insights and challenges in the dynamic environment of sequential decision-making tasks.
3. Reinforcement Unlearning
3.1. Problem Statement and Threat Model
Problem Definition.  The definition of “forgetting” is application-dependent, leading to different desiderata and priorities in various scenarios (Kurmanji et al., 2023). For example, in a privacy-centric application, the main goal of unlearning user data is to ensure that the unlearned model has no exposure to the data, and a successful membership inference attack would reveal that the data is not in the training set for the unlearned model. Conversely, in a bias-removing application, the aim of unlearning is to prevent the unlearned model from predicting the assigned labels of the forgotten data, as these labels may indicate unintended and biased behavior.
The objective of reinforcement unlearning is to eliminate the influence of a specific environment on the agent, i.e., “forgetting an environment”. We define “forgetting an environment” as equivalent to “performing deterioratively in that environment”. This definition aligns with common sense. For example, when we have thoroughly explored a place and are highly familiar with it, we can efficiently find things within it, resulting in high performance. Conversely, when we have forgotten a place, our ability to locate things diminishes, leading to deteriorative performance.
It is worth noting that we cannot simply adopt the concept of “forgetting” from conventional machine unlearning, which often involves retraining the model from scratch without the revoked data. This is because reinforcement unlearning operates within a distinct learning paradigm, such as sequential decision making and dynamic learning. For instance, even if we were to retrain an agent from scratch without including a specific unlearning environment, there’s a possibility that the agent still performs well in that environment. However, this positive performance could be exploited by adversaries to deduce critical information about that environment. This clearly contradicts the fundamental objective of safeguarding the environment owner’s privacy in reinforcement unlearning.
@@ 3.1. Problem Statement and Threat Model
(3) $\min_{\pi^{\prime}}||Q_{\pi^{\prime}}(s)||_{\infty},$
@@
(4) $\min_{\pi^{\prime}}||Q_{\pi^{\prime}}(s)-Q_{\pi}(s)||_{\infty},$
where $s\notin\mathcal{S}_{u}$. We assume that the owner of the trained RL model can access $\mathcal{M}_{u}$ and gather trajectories within $\mathcal{M}_{u}$. A trajectory $\tau$ is denoted as a sequence of state-action pairs: $\tau=((s_{1},a_{1}),\ldots,(s_{k},a_{k}))$, where $k$ represents the length of the trajectory.
The assumption of the model owner having access to the unlearning environment $\mathcal{M}_{u}$ is reasonable, as $\mathcal{M}_{u}$ is an integral part of the training data used by the model owner to train the agent. When an unlearning request is initiated, the model owner employs the proposed unlearning methods, leveraging the unlearning environment $\mathcal{M}_{u}$, to execute the unlearning process. Subsequently, the model owner physically removes $\mathcal{M}_{u}$. This approach ensures that the unlearning is performed using the relevant environment data under the ownership of the user, maintaining a practical and privacy-conscious procedure.
@@
It is essential to note that the environment inference attack differs significantly from conventional membership inference attacks in machine learning. Traditional membership inference attacks typically involve point-level inference, where the focus is on deducing information about an individual sample. In contrast, the environment inference attack operates at the object level, concentrating on the inference of features characterizing an entire environment that encompasses a substantial number of samples. This distinction underscores the unique nature of the environment inference attack, as it extends its scope beyond individual data points to involve the broader context of environments, introducing a new dimension to the evaluation methodology for reinforcement unlearning.
@@
Methods Overview.  Both decremental reinforcement learning-based method and the poisoning-based method share the common aim of intentionally degrading the agent’s performance within the unlearning environment while preserving its performance in other environments. However, they employ distinct strategies to achieve this outcome. The decremental reinforcement learning-based method involves updating the agent by minimizing its reward specifically in the unlearning environment. This is achieved through iterative adjustments to the agent’s policy, aiming to reduce its effectiveness within the unlearning environment. In contrast, the environment poisoning-based method focuses on modifying the unlearning environment itself. This method involves introducing deliberate changes to the state transition function of the environment and subsequently updating the agent in this modified environment. The intention is to disrupt the agent’s learned behavior in the unlearning environment.
3.2. Decremental RL-based Method

### v5

@@ Abstract
Machine unlearning refers to the process of mitigating the influence of specific training data on machine learning models based on removal requests from data owners. However, one important area that has been largely overlooked in the research of unlearning is reinforcement learning. Reinforcement learning focuses on training an agent to make optimal decisions within an environment to maximize its cumulative rewards. During the training, the agent tends to memorize the features of the environment, which raises a significant concern about privacy. As per data protection regulations, the owner of the environment holds the right to revoke access to the agent’s training data, thus necessitating the development of a novel research field, termed reinforcement unlearning. Reinforcement unlearning focuses on revoking entire environments rather than individual data samples. This unique characteristic presents three distinct challenges: 1) how to propose unlearning schemes for environments; 2) how to avoid degrading the agent’s performance in remaining environments; and 3) how to evaluate the effectiveness of unlearning. To tackle these challenges, we propose two reinforcement unlearning methods. The first method is based on decremental reinforcement learning, which aims to erase the agent’s previously acquired knowledge gradually. The second method leverages environment poisoning attacks, which encourage the agent to learn new, albeit incorrect, knowledge to remove the unlearning environment. Particularly, to tackle the third challenge, we introduce the concept of “environment inference” to evaluate the unlearning outcomes. The source code is available at https://github.com/cp-lab-uts/Reinforcement-Unlearning.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 23 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.23080 www.ndss-symposium.org
I Introduction
@@
Machine learning relies on the acquisition of vast amounts of data. To safeguard the data privacy of individual users, data protection regulations, e.g., the General Data Protection Regulation [11], empower users to request the removal of their data. It is imperative for model owners to adhere to users’ requests by removing revoked data from their datasets and ensuring that any influence these revoked data may have on the model is eliminated. This process is referred to as machine unlearning [3, 2].
@@
While significant progress has been made in conventional machine unlearning [2, 13, 48], one area that remains an unfilled gap for unlearning is reinforcement learning (RL). RL is an essential research field in machine learning due to its ability to address complex decision-making problems in dynamic environments [42]. In RL, the primary objective is to train an intelligent entity, known as an agent, to interact with the environment through a specific policy. This policy guides its actions based on the current state. With each action taken, the agent receives a reward and updates its state, creating an experience sample used to update its policy. The aim of the agent is to learn an optimal policy that maximizes its cumulative rewards over time.
@@
However, in the course of RL, agents tend to memorize features of their environments, raising security concerns. Consider an agent designed for providing navigation guidance through real-time data from Google Maps. During its training, the agent learns from a dynamic environment using Google Street Views for its photographic content [25]. However, privacy issues may arise when the agent inadvertently learns sensitive information, e.g., the locations of restricted areas.
Another example arises from RL-based recommendation systems. Major platforms like YouTube [4], Netflix [36], and Amazon [28] have successfully implemented RL to enhance the quality of their recommendations, providing users with more personalized services. In these systems, each user’s interaction history and preferences represent a unique environment. The system, i.e., the agent, learns to make personalized recommendations by adapting its strategy to each user’s specifics, which often contain sensitive information, such as purchasing habits. When users opt out or request data deletion, the recommendation system must comply by forgetting users’ specifics, giving rise to the novel field of reinforcement unlearning.
@@
Reinforcement unlearning is also distinct from privacy-preserving RL [39, 10]. Reinforcement unlearning aims to selectively erase learned knowledge from the agent’s memory, ensuring the privacy of environment owners, while privacy-preserving RL focuses on preserving the agent’s personal information. In essence, reinforcement unlearning presents three specific challenges.
  • How can we unlearn an environment from the agent’s policy? In machine unlearning, a data owner can specify which data samples should be removed. However, in reinforcement unlearning, the environment owner cannot access the experience samples, as these samples are managed by the agent. Thus, the key challenge lies in effectively associating the environment that needs to be unlearned with the corresponding experience samples.
  • How can we prevent a degradation in the agent’s performance after unlearning? In conventional machine unlearning, removing samples may lead to a decrease of performance. It is more challenging in reinforcement unlearning as unlearning an environment requires forgetting a significant number of experience samples.
  • How can we evaluate the effectiveness of reinforcement unlearning? In machine unlearning, one commonly used evaluation is using membership inference  [34] to assess if the model has forgotten the revoked data. However, this methodology cannot be directly applied to reinforcement unlearning, as the environment owner cannot specify which samples should be unlearned.
To address these challenges, we propose two distinct unlearning methods: decremental reinforcement learning and environment poisoning. Decremental reinforcement learning involves deliberately erasing an agent’s learned knowledge about a specific environment. This method can be applied to scenarios where certain environments become obsolete or need to be forgotten due to privacy concerns. Environment poisoning-based method aims to create poisoning experience samples by modifying the unlearning environment. This method ensures that the agent’s performance in other environments remains unaffected, eliminating any negative impact on its overall capabilities. This method finds application in situations where attacks or misinformation may be present. Both methods enable an agent to unlearn specific environments while maintaining its performance in others, thereby tackling the first two challenges. To tackle the third challenge, we utilize environment inference to infer an agent’s training environment by observing its behavior. If the inference result after unlearning shows a substantial degradation compared to the result before unlearning, it is indicative that the agent has effectively unlearned that environment.
@@
  • Reinforcement unlearning exposes an impactful vulnerability of RL – the risk of exposing the privacy of the environment owner. This vulnerability can disclose sensitive information about the environment owner’s preferences and intentions. We introduce two innovative reinforcement unlearning methods: decremental RL-based and environment poisoning-based approaches to tackle this vulnerability.
  • To confirm the unlearning results, we introduce a novel evaluation approach, “environment inference”. By visualizing the unlearning results, this approach provides an intuitive means of measuring the efficacy of unlearning techniques.
II Preliminaries
The notations used in this paper are listed in Table I.
TABLE I: Summary of Notations
Notations                           Description
$\mathcal{M}$ A learning environment/task
$\mathcal{S}$ A state set, $\mathcal{S}=\{s_{1},...,s_{n}\}$
$\mathcal{A}$ An action set, $\mathcal{A}=\{a_{1},...,a_{m}\}$
$\mathcal{T}$ A transition function
$r$ A reward function
$\gamma$ The discount factor
$\pi$ A policy learned by an agent
$Q(s,a)$ The value of a $Q$-function by taking action $a$ in state $s$
$e$ An experience sample
$B$ A batch of experience samples for agent learning
$\mathcal{B}$ A batch of experience samples for poisoning strategy update
$\tau$ A trajectory consisting of a sequence of state-action pairs
$||\mathbf{x}||_{\infty}$ The maximum dimension in a vector $\mathbf{x}$
Reinforcement Learning.  In RL, a learning environment is formulated by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},r\rangle$ [26, 35]. $\mathcal{S}$ and $\mathcal{A}$ denote the state and action sets, respectively, while $\mathcal{T}$ represents the transition function, and $r$ represents the reward function. At each time step $t$, the agent, given the current environmental state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$ based on its policy $\pi(s_{t},a_{t})$. This action causes a transition in the environment from state $s_{t}$ to $s_{t+1}$ according to the transition function: $\mathcal{T}(s_{t+1}|s_{t},a_{t})$. The agent then receives a reward $r_{t}(s_{t},a_{t})$, along with the next state $s_{t+1}$. This tuple of information, denoted as $(s_{t},a_{t},r_{t}(s_{t},a_{t}),s_{t+1})$, is collected by the agent as an experience sample utilized to update its policy $\pi$. Typically, the policy $\pi$ is implemented using a Q-function: $Q(s,a)$, estimating the accumulated reward the agent will attain in state $s$ by taking action $a$. Formally, the Q-function is defined as:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}[\sum^{\infty}_{i=1}\gamma^{i}\cdot r(s_{i},a_{i})|s_{i}=s,a_{i}=a],$ (1)
@@ II Preliminaries
In deep RL, a neural network is employed to approximate the Q-function [27], denoted as $Q(s,a;\theta)$, where $\theta$ represents the weights of the neural network. The neural network takes the state $s$ as input and produces a vector of Q-values as output, with each Q-value corresponding to an action $a$. To learn the optimal values of $Q(s,a;\theta)$, the weights $\theta$ are updated using a mean squared error loss function $\mathcal{L}(\theta)$.
$\mathcal{L}=\frac{1}{|B|}\sum_{e\in B}[(r(s_{t},a_{t})+\gamma\max_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta)$ (2)
$-Q(s_{t},a_{t};\theta))^{2}],$
@@
Machine Unlearning.  Machine unlearning aims to erase the impact of certain data samples on a trained model’s behavior [3]. A straightforward unlearning method is removing the revoked data and retraining the model from scratch. Formally, given a dataset $D$, an unlearning set $D_{u}$ and a learning algorithm $A$, the objective of machine unlearning is to guarantee that the unlearned model $A(D/D_{u})$ performs as if it had never seen the unlearning set $D_{u}$.
III Reinforcement Unlearning
III-A Problem Statement and Threat Model
Problem Definition.  The definition of “forgetting” is application-dependent [20]. For example, in a privacy-centric application, the main goal of unlearning user data is to ensure that the unlearned model has no exposure to the data, and a successful membership inference would reveal that the data is not in the training set for the unlearned model. Conversely, in a bias-removing application, the aim of unlearning is to prevent the unlearned model from predicting the assigned labels of the forgotten data, as these labels may indicate unintended and biased behavior.
The objective of reinforcement unlearning is to eliminate the influence of a specific environment on the agent, i.e., “forgetting an environment”. We define “forgetting an environment” as equivalent to “performing deterioratively in that environment”. This definition aligns with common sense. For example, when we have thoroughly explored a place and are highly familiar with it, we can efficiently find things within it, resulting in high performance. Conversely, when we have forgotten a place, our ability to locate things diminishes, leading to deteriorative performance. This definition may not align with the conventional machine unlearning approach, which usually focuses on ensuring the model’s performance matches that of one retrained from scratch. This difference arises because reinforcement unlearning operates within a distinct learning paradigm, characterized by sequential decision-making and dynamic learning. Our definition is motivated by the need to address privacy concerns and mitigate the impact of sensitive or incorrect information within the environment.
@@ III-A Problem Statement and Threat Model
$\min_{\pi^{\prime}}||Q_{\pi^{\prime}}(s)||_{\infty},$ (3)
@@
$\min_{\pi^{\prime}}||Q_{\pi^{\prime}}(s)-Q_{\pi}(s)||_{\infty},$ (4)
where $s\notin\mathcal{S}_{u}$. Note that if the agent needs to regain knowledge of previously unlearned environments, it can treat them as new environments. This process involves employing an incremental learning approach, where the model is fine-tuned by focusing specifically on the distinct features of these environments.
We assume that the owner of the trained RL model can access $\mathcal{M}_{u}$ and gather trajectories within $\mathcal{M}_{u}$. A trajectory $\tau$ is denoted as a sequence of state-action pairs: $\tau=((s_{1},a_{1}),\ldots,(s_{k},a_{k}))$, where $k$ is the length of the trajectory. In scenarios where environments become inaccessible after training, the agent stores the experience samples collected during its training and labels them based on their respective environments. These labeled samples can be used to implement the proposed unlearning methods by separating the samples into those from the unlearning environment and those from the remaining environments.
Note that our work can also be applied to unlearning in a single environment, where the agent is trained within one environment and aims to selectively forget specific aspects of that environment. Detailed discussion and experimental results regarding this adaptation are included in Appendix I.
@@
Environment inference attack differs significantly from conventional membership inference attacks in machine learning. Membership inference attacks typically involve point-level inference, where the focus is on deducing information about an individual sample. In contrast, the environment inference attack operates at the object level, concentrating on the inference of features characterizing an entire environment that encompasses a substantial number of samples.
@@
Methods Overview.  Both decremental RL-based and the poisoning-based methods share the common aim of intentionally degrading the agent’s performance within the unlearning environment while preserving its performance in other environments. However, they employ distinct strategies to achieve this outcome. The decremental RL-based method involves updating the agent by minimizing its reward specifically in the unlearning environment. This is achieved through iterative adjustments to the agent’s policy, aiming to reduce its effectiveness within the unlearning environment. In contrast, the environment poisoning-based method focuses on modifying the unlearning environment itself. This method involves introducing deliberate changes to the state transition function of the environment and subsequently updating the agent in this modified environment. The intention is to disrupt the agent’s learned behavior in the unlearning environment.
III-B Decremental RL-based Method